OpenStudy (carissa15):

Please help. I have a calculus question about the secant of a parabola y=x^2 that passes through points A=(2,4) and A'=(2+delta x, 4+ delta y) , find the slope of this secant. I am then given a) delta x =0.01 and b) delta x = E. To answer this do I just sub in the a) and b) values of x and solve the equation?

OpenStudy (michele_laino):

hint: the slope m of your secan t, is: \[m = \frac{{{\delta _y}}}{{{\delta _x}}}\]

OpenStudy (michele_laino):

furthermore, we can write this: \[\Large 4 + {\delta _y} = {\left( {2 + {\delta _x}} \right)^2}\]

OpenStudy (carissa15):

you then solve for delta y? As values for delta x are already given?

OpenStudy (michele_laino):

from my last equation, we have: \[\Large 4 + {\delta _y} = 4 + {\left( {{\delta _x}} \right)^2} + 4{\delta _x}\] so we can simplify as below: \[\Large \frac{{{\delta _y}}}{{{\delta _x}}} = {\delta _x} + 4\]

OpenStudy (carissa15):

hmm

ganeshie8 (ganeshie8):